Project

General

Profile

2018-04-03-MEMOMIPO

Joint MICE Experiment Management Office and MICE International Project Office

3rd April 2018; 15:30 BST
RAL, R1 2nd Floor East, CR26

Present: KL, CW, PS, DR, PG, ABr, JB
Apologies: AK, SB, CR

  1. Introduction: KL
    • IPAC -- VC; Equipment to return.
  2. Notes of previous meeting and status of actions: All
  3. Hall status, stability of systems, planning for decommissioning and making safe to be complete by 31Mar18: CW
    • Field mapping is complete. Spectrometer solenoids have been warmed up. Power supplies are now locked out.
    • Hall is now on controlled to reduce risk that equipment is removed.
    • New crates will be fabricated for the EMR and perhaps also for ToF0/1.
    • PS reminds us that the ADCs belong to Italy while the TDCs belong to the UK.
  4. Plans for equipment repatriation: CW
    • LiH needs to be inspected at RAL before it is sent. This will be organised.
    • Spreadsheet from ABr contains list of equipment that should be returned to the US. CW will add this list to the document summarising all equipment to be returned.
    • DR Reminds us that some computers in the RR that should be returned to GVA and the US. Possible that US inventry includes a computed in the Tracker Lab and in the Hall.
  5. Software and computing status: DR
    • To include:
      • Processing for scattering paper; TOF geometry update; tracker and ToF s/w review uptdate; S/w&C decommissioning and staff plan.
      • Data processing for the scattering paper is underway. Similarly, processing for emittance paper is underway.
      • Tracker s/w review close out will be in parallel to the analysis w/s in Ggo.
      • The ToF calibration s/w update now includes the development of a pixel-pairing algorithm to improve the calibration coverage. The is underway and DR expects a status report next.
      • Data/certificate-management personnel are being sought in two areas: local support at RAL and infrequent support that can be off-site. It is conceivable that off-site support will be provided by Brunel. Kyberd, Rajaram and Nebrensky will meet this week to work out how to cover the topics. Goal will be to close out with KL by the end of the week.
      • MAUS paper is being editted by DR. Next draft will be circulated by the end of the week.
  6. Paper status and plans: CR
    • To include:
      • Status of preperations for IPAC18.
      • CR not available so no report this time. KL reports for CR:
        • Emittance paper:
          • Authors have profuced hat they consider to be an essentially final draft of the MICE note. A referees meeting is planed for 09Apr18.
        • Scattering paper:
          • The most significant issue to be addressed is the asymmetry in the empty-channel scattering distribution (and the difference between the x- and y-scattering distributions. In addition, the convergence of the unfolding technique needs to be addressed and JC is arguing for a change in the acceptance-correction procedure.
      • IPAC18:
        • Scattering update: it had been hoped that final LiH scattering distributions would be presented. This analysis is not quite complete (see above).
        • Emittance/amplitude evolution: good progress is being made by CR. Remaining issues include a consistent/complete estimation of the systematic errors. Await presentation at Thursday's VC.
        • Wedge analysis: it is unclear whether TM has brought the analysis to the stage where she will be able to present it to the collaboration for ask that the results be made public. KL will clarify.
        • Diffuser analysis: MU has been unwell and therefore has been forced to withdraw her abstract.
  7. AOB
    • Close-out review: draft of documentation for will be made available for the EB on Friday 06Apr18.

Summary of actions:

None.

Back to MEMO:

Back

Back to MIPO:

Back

Documents

2018-04-03-MICE-collaborator-equipment.pdf (328 KB) Long, Kenneth, 03 April 2018 09:36